The formula for the surface area of a sphere is A=4pier^2.A square sticker of side x is placed on a ball of radius 3x.What's the surface area of the sphere not covered by the sticker?Write your answer in factored form.
Answer by edjones(8007)

A=4*pi*r^2 (sphere)
=4*pi*(3x)^2
=4*pi*9x^2
=pi*36x^2 or 113.01x^2
.
A=x^2 (sticker)
.
pi*36x^2-x^2 or 112.01x^2 Surface area of the sphere not covered by the sticker.
.
Ed